XYZ Company purchased a new machine on January 1, 2021 for $80,000.
The machine was assigned a 10-year life and a $5,000 residual value.
XYZ Company will depreciate the machine using the double-declining
balance depreciation method.

Calculate the amount of accumulated depreciation related to the new
machine that would appear in XYZ Company's December 31, 2024 balance
sheet.
Answer #1

Answer

  • Working

A

Cost

$            80,000.00

B

Residual Value

$              5,000.00

C=A - B

Depreciable base

$            75,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation

$              7,500.00

F=E/C

SLM Rate

10.00%

G=F x 2

DDB Rate

20.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

2021

$              80,000.00

20.00%

$         16,000.00

$             64,000.00

$        16,000.00

2022

$              64,000.00

20.00%

$         12,800.00

$             51,200.00

$        28,800.00

2023

$              51,200.00

20.00%

$         10,240.00

$             40,960.00

$        39,040.00

2024

$              40,960.00

20.00%

$            8,192.00

$             32,768.00

$        47,232.00

  • Correct Answer:
    Accumulated Depreciation on 31 Dec 2024 = $ 47,232